OpenStudy (anonymous):

A group of 11 people are going to the movies one Saturday. Adult tickets cost $13.50 each and child tickets cost $8.25 each. If the total cost for admission to the movies is $111.75, how many children attended the movie?

OpenStudy (campbell_st):

let a = adult and c = child and you know a + c = 11 number of people 13.5a + 8.25c = 111.75 cost take the number of people equation and make a the subject a = 11 - c now substitute it into the cost equation 13.5(11 - c) + 8.25c =111.75 now simplify then solve for c

OpenStudy (anonymous):

ok i got 6

OpenStudy (anonymous):

is that right?

OpenStudy (campbell_st):

well I got 7

OpenStudy (anonymous):

ohh ok i see what i did wrong